Grade 12th passAlgebra

Find the measure of each exterior angle of a regular polygon of 15 sides

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Sum of angles a regular polygon having side n = (n-2)×180°
Sum of angles a regular polygon having side 15 = (15-2)×180°
= 13×180° = 2340°
Each interior angle = 2340/15 = 156°
Each exterior angle = 180° - 156° = 24°
Or,
Each exterior angle = sum of exterior angles/Number of angles =
360/15 = 24°

Thanks